# Rate, Explain and Cite (REC): Enhanced Explanation and Attribution in Automatic Evaluation by Large Language Models

### v2

@@ Abstract
LLMs have demonstrated impressive proficiency in generating coherent and high-quality text, making them valuable across a range of text-generation tasks. However, rigorous evaluation of this generated content is crucial, as ensuring its quality remains a significant challenge due to persistent issues such as factual inaccuracies and hallucination. This paper introduces three fine-tuned general-purpose LLM autoevaluators, REC-8B, REC-12B and REC-70B, specifically designed to evaluate generated text across several dimensions: faithfulness, instruction following, coherence, and completeness. These models not only provide ratings for these metrics but also offer detailed explanation and verifiable citation, thereby enhancing trust in the content. Moreover, the models support various citation modes, accommodating different requirements for latency and granularity. Extensive evaluations on diverse benchmarks demonstrate that our general-purpose LLM auto-evaluator, REC-70B, outperforms state-of-the-art LLMs, excelling in content evaluation by delivering better quality explanation and citation with minimal bias. It achieves Rank #1  [^1] As of Feb 15th, 2025as a generative model on the RewardBench leaderboard [^2] https://huggingface.co/spaces/allenai/reward-bench under the model name TextEval-Llama3.1-70B. Our REC dataset and models are available at https://github.com/adelaidehsu/REC.
Rate, Explain and Cite (REC): Enhanced Explanation and Attribution in Automatic Evaluation by Large Language Models
Aliyah R. Hsu1,2thanks: Work done during an internship at Salesforce., James Zhu1, Zhichao Wang1, Bin Bi1, Shubham Mehrotra1, Shiva K. Pentyala1, Katherine Tan1, Xiang-Bo Mao1, Roshanak Omrani1, Sougata Chaudhuri1, Regunathan Radhakrishnan1, Sitaram Asur1, Claire Na Cheng1, Bin Yu2 1Salesforce AI Platform, 2UC Berkeley Correspondence: aliyahhsu@berkeley.edu, james.zhu@salesforce.com
@@ 1 Introduction
Figure 1: Illustration of the REC framework for content quality evaluation. The autoevaluator takes in context (a task prompt from another LLM), generation from another LLM, and a user-specified evaluation metric (i.e., completeness). The autoevaluator outputs rating, explanation with citation according to a user-specified citation mode (i.e., inline with context snippet). Citation are extracted verbatim from context as underlined. For details of the alternative citation mode and the evaluation prompt, see Appendix A.
@@
Our models are the first to enable citation in both automatic evaluation (i.e., content quality citation) and general task output (i.e., general RAG citation) with scalability and efficiency, while prior work often focus on providing solely rating and explanation in automatic evaluation (Zhu et al., 2023; Li et al., 2023a; Wang et al., 2023b; Vu et al., 2024a), or require iterative prompting (Sun et al., 2023; Cao and Wang, 2024; Ye et al., 2023) or human annotation (Malaviya et al., 2024) for general citation generation. Our contributions include the following:
  • A novel general-purpose LLM autoevaluator that comes in three sizes: 8B, 12B and 70B. Our LLM autoevaluators can generate improved quality Rating, Explanation and Citation (REC), with little to no trade-off in general instruction task performance.  [^3] All evaluations are limited to English datasets.
@@
We provide extensive evaluations of our REC models on diverse benchmarks evaluating LLMs’ (1) RAG citation capability: ALCE (Gao et al., 2023), and ExpertQA (Malaviya et al., 2024); (2) content quality citation capability: human evaluation on ABCD summarization (Chen et al., 2021); (3) general capabilities: RewardBench (Lambert et al., 2024), and LLM-AggreFact (Tang et al., 2024a); (4) cognitive bias: CoBBLEr (Koo et al., 2024), against 7 state-of-the-art (SOTA) LLMs, such as GPTs and Claude. We show that our REC models outperform state-of-the-art baselines on these benchmarks, and that the baselines often struggle to provide accurate citation, undermining their reliability. Specifically, our REC-70B model achieves Rank #1  [^5] As of Feb 15th, 2025 as a generative model on the RewardBench leaderboard under the model name TextEval-Llama3.1-70B. The model outperforms the second-ranked model in three dimensions: Chat Hard, Safety, Reasoning, which are key capabilities required to generate accurate explanation and citation outputs.
@@ 2 REC Model Citation Capabilities
We illustrate the content quality citation task in Figure 1, where the autoevaluator performs content quality evaluation on another LLM’s generation according to a specified metric, and provides rating and explanation with citation for verifiability as an output. When not in an evaluation setting, our models can perform general RAG citation, where we cite from retrieved articles and tag the citation with an answer generated by another LLM, as illustrated in Figure 2.
@@ 3 Related Work
Providing a reference to the attributable source of information (Liu et al., 2023a; Gao et al., 2023) has been proposed as one of the approaches to mitigate hallucination and improve the factuality of LLM generation. Such work fall into the general RAG citation setting defined in our work, where the aim is to generate verifiable content through citation of provided articles. Initial efforts fine-tune LLMs using human-written examples (Nakano et al., 2021) or machine-generated examples verified by humans (Menick et al., 2022), but their privately maintained training data limits further research. With the advent of more capable LLMs (OpenAI, 2023; Jiang et al., 2023b), most existing work rely on zero-shot prompting or few-shot prompting to cite articles during generation (Kamalloo et al., 2023; Gao et al., 2023; Liu et al., 2023a), although the quality of their generated citation leaves much room for improvement (Malaviya et al., 2024). Other work utilize an additional natural language inference (NLI) model to add citation (Gao et al., 2022; Chen et al., 2023). To improve upon prior work, we propose a fine-tuned LLM that can generate better-grounded responses supported with citation of various modes, with the fine-tuning dataset released to the public to facilitate future research and no additional NLI models needed.
General-purpose Autoevaluators
Collecting human annotations to evaluate LLM is not only costly and time consuming, but hard to replicate (Ouyang et al., 2022; Zheng et al., 2023; Chiang and Lee, 2023), and as a result, LLMs become a natural automated proxy to evaluate LLM capabilities on various benchmarks (Bai et al., 2024; Bubeck et al., 2023; Chiang et al., 2023; Fu et al., 2023; Liu et al., 2023b; Wang et al., 2023a). Existing LLM autoevaluators often judge LLM outputs by expressing “preference” over outputs from a reference model (Dubois et al., 2024; Li et al., 2023b; Yuchen Lin et al., 2024), or by providing rating and explanation according to a user-defined metric as a direct assessment (Li et al., 2023a; Wang et al., 2023b, 2024b). Closer to our work, Jiang et al. (2024) and Xu et al. (2023) fine-tune LLMs to generate rating, explanation, and a detailed analysis to pinpoint errors in a response evaluated. Unlike prior work, our fine-tuned model is a more generalizable autoevaluator since it provides content quality citation for both good and bad evaluated responses, in addition to rating and explanation. This not only allows users to be able to diagnose where errors are from but provides supporting evidence when a generation is good, which are both important to establish trust in a model.
@@ 4 Data Collection
To fine-tune our general-purpose LLM autoevaluator, we meticulously collect a mixture of data encompassing a broad spectrum of LLM capabilities (Section 4.1), and we denote our collected dataset as REC-Data. Due to a lack of publicly available content quality citation datasets, we leverage synthetic data generation using an LLM to curate such data (Section 4.2) to facilitate the research community. We perform rule-based automatic quality check followed by a unified task formatting to post-process our data for instruction fine-tuning (Section 4.3).
@@ 4.1 Task Types
To enhance the explainability in the automatic evaluation of our LLM autoevaluator, while maintaining its general instruction-following capability, we gather datasets from a diverse range of task types (See detailed REC-Data distribution in Appendix C). These tasks represent essential capabilities that an advanced autoevaluator LLM should possess:
@@ 5 Model Training
Our general-purpose LLM autoevaluator is available in three sizes, REC-8B, REC-12B, and REC-70B. REC-12B is instruction fine-tuned from Mistral-Nemo [^7] https://mistral.ai/news/mistral-nemo, and REC-8B and REC-70B are instruction fine-tuned from Llama-3.1-8B and Llama-3.1-70B Team (2024), respectively. We trained REC-8B and REC-12B to assess the generalizability of the REC approach to different model architectures in similar model sizes. We adopted supervised fine-tuning (SFT) to optimize the models. To accommodate GPU memory constraints: we filtered examples where the combined length of a prompt and a response exceeded 6,144 tokens. We used Low-Rank Adaptation (LoRA) during fine-tuning Hu et al. (2021), with a rank of $r=256$ for REC-8B and REC-12B, and $r=64$ for REC-70B. We applied 4-bit quantization for REC-70B during initialization. All models were trained using a learning rate of $1\times 10^{-4}$, with data shuffled and for a single epoch. REC-8B and REC-12B were trained with a batch size of 2 and a gradient accumulation factor of 8, whereas for REC-70B we used a batch size of 1 with the same gradient accumulation factor. All models were trained on eight NVIDIA H100 GPUs, each with 80GB of memory. The total training time was approximately 4.5 hours for REC-8B and REC-12B, and 5.5 hours for REC-70B.
@@ 6 Evaluation
In this section, we evaluate our REC models on diverse benchmarks assessing LLMs’ (1) RAG citation capability: ALCE (Gao et al., 2023), and ExpertQA (Malaviya et al., 2024); (2) content quality citation capability: human evaluation on ABCD summarization (Chen et al., 2021); (3) general capabilities: RewardBench (Lambert et al., 2024), and LLM-AggreFact (Tang et al., 2024a); (4) cognitive bias: CoBBLEr (Koo et al., 2024), each of which is an important measure to a general-purpose LLM autoevaluator. We compare our results against 7 SOTA LLMs, including Misral-7B (Mistral-7B-Instruct-v0.2(Jiang et al., 2023a), Mistral-Nemo (Mistral-Nemo-Instruct-2407), Llama-3.1-70B (Grattafiori et al., 2024), Claude-3-Opus (claude-3-opus-20240229), GPT-3.5 (gpt-3.5-turbo), GPT-4 (gpt-4-turbo(OpenAI et al., 2024), and GPT-4o (gpt-4o).
@@ 6.1 RAG citation & Correctness
ExpertQA Malaviya et al. (2024) is designed to evaluate LLMs’ ability to generate accurate and well-attributed responses in technical and high-stakes fields, such as medicine and law. It was developed by involving experts from 32 different fields, who contributed 2,177 domain-specific questions based on their knowledge. In total, 484 participants helped curate these questions. LLMs were then used to generate responses to the questions, followed by human experts evaluating the quality of these responses on several criteria, including factual correctness, completeness of attribution, source reliability, and informativeness. We adopt the metrics used in its paper: AutoAIS Gao et al. (2022), which is similar to citation recall, and FActscore Min et al. (2023) which measures the percentage of generated claims that are factual.
@@ 6.2 Rating & Explanation & Citation
To evaluate our LLM autoevaluator’s citation capability more comprehensively, besides examining its general RAG citation capability on QA tasks as reported in Section 6.1, we evaluate the content quality citation capability on a summarization task, the ABCD dataset Chen et al. (2021), in this section.
@@
The ABCD dataset Chen et al. (2021) contains customer support conversation transcripts. We generate summaries by prompting an open-source LLM, Mistral-7B (Jiang et al., 2023b), with: “Summarize knowledge from transcripts after they’ve ended, including the customer issue and resolution.” to form (dialogue, summary) pairs for evaluation. LLM autoevaluators then take as input the summarization task prompt, and a (dialogue, summary) pair, to generate rating, explanation, and citation according to the metrics defined in Section 1. Note that we use a held-out set of 50 examples in this experiment so there exists no data overlap with the ABCD subsets used during training.
@@
Unlike the general RAG citation benchmarks evaluated in Section 6.1, which contain human annotated citation groundtruth, there exists no well-established benchmarks to evaluate content quality citation. Hence, we seek human annotations to serve as a reference standard in this task. The machine outputs were labeled by 12 machine learning experts, each of whom has a graduate degree in computer science or related fields. They were asked to complete three tasks: rating correctness evaluation, explanation correctness evaluation, and citation. For rating and explanation evaluation, the annotators were asked to provide “Yes/No” labels to determine the correctness of the responses generated by the models. For citation evaluation, they were asked to to provide manually-written citation for the claims in the generated explanation to serve as the groudtruth citation. (Details of the human labeling guideline can be found in Appendix F). Human evaluation on the Coherence metric was not performed because it is conceptually challenging to define what should be cited in a coherent response (e.g., citing an entire paragraph adds little interpretive value). Citation for coherence are more meaningful in negative cases where parts of the response are incoherent. To ensure a balanced evaluation with both positive and negative cases across metrics, we decided to exclude Coherence from the ABCD summarization content quality task.
@@
We assign two labelers for each model and find the average inter-rater agreement is around 95%. We report the average accuracy for rating and explanation across the two sets of labels, and we compute F1 score between machine-generated citation and the intersection of the two sets of manual citation. From Table 3, we see that most LLM autoevaluators obtain high rating and explanation accuracies, indicating that rating in “Yes/No” and free-form explanation generation are relatively easy tasks for the LLMs. For citation, our REC models perform the best in both completeness and faithfulness metrics, while in instruction-following, they are ranked second to llama-3.1-70B and Claude3-Opus. Specifically, for the instruction-following metric, we find REC models often cite slightly more than necessary sentences that appear to cover a conversation between a customer and an agent. In contrast, Llama-70B and Claude3 seem to be able to selectively cite the customer’s complaint and the agent’s resolution to reflect the instruction to focus on these two elements in the summary.
@@
Table 4: General LLM capabilities evaluation on RewardBench. Scores for all the compared LLMs are obtained from the RewardBench Leaderboard Lambert et al. (2024).

### v3

@@ Abstract
LLMs have demonstrated impressive proficiency in generating coherent and high-quality text, making them valuable across a range of text-generation tasks. However, rigorous evaluation of this generated content is crucial, as ensuring its quality remains a significant challenge due to persistent issues such as factual inaccuracies and hallucination. This paper introduces three fine-tuned general-purpose LLM auto-evaluators, REC-8B, REC-12B and REC-70B, specifically designed to evaluate generated text across several dimensions: faithfulness, instruction following, coherence, and completeness. These models not only provide ratings for these metrics but also offer detailed explanation and verifiable citation, thereby enhancing trust in the content. Moreover, the models support various citation modes, accommodating different requirements for latency and granularity. Extensive evaluations on diverse benchmarks demonstrate that our general-purpose LLM auto-evaluator, REC-70B, outperforms state-of-the-art LLMs, excelling in content evaluation by delivering better quality explanation and citation with minimal bias. Our REC dataset and models are available at https://github.com/adelaidehsu/REC.
Explainable AI, Trustworthy AI, Large Language Models
@@ 1 Introduction
Figure 1: Illustration of the REC framework for content quality evaluation. The auto-evaluator takes in context (a task prompt from another LLM), generation from another LLM, and a user-specified evaluation metric (i.e., completeness). The auto-evaluator outputs rating, explanation with citation according to a user-specified citation mode (i.e., inline with context snippet). Citation are extracted verbatim from context as underlined. For details of the alternative citation mode and the evaluation prompt, see Appendix A.
@@
Our models are the first to enable citation in both automatic evaluation (i.e., content quality citation) and general task output (i.e., general RAG citation) with scalability and efficiency, while prior work often focus on providing solely rating and explanation in automatic evaluation (Zhu et al., 2023; Li et al., 2023a; Wang et al., 2023b; Vu et al., 2024a), or require iterative prompting (Sun et al., 2023; Cao & Wang, 2024; Ye et al., 2023) or human annotation (Malaviya et al., 2024) for general citation generation. Our contributions include the following:
  • A novel general-purpose LLM auto-evaluator that comes in three sizes: 8B, 12B and 70B. Our LLM auto-evaluators can generate improved quality Rating, Explanation and Citation (REC), with little to no trade-off in general instruction task performance.  [^1] All evaluations are limited to English datasets.
@@
We provide extensive evaluations of our REC models on diverse benchmarks evaluating LLMs’ (1) RAG citation capability: ALCE (Gao et al., 2023), and ExpertQA (Malaviya et al., 2024); (2) content quality citation capability: human evaluation on ABCD summarization (Chen et al., 2021); (3) general capabilities: RewardBench (Lambert et al., 2024), and LLM-AggreFact (Tang et al., 2024a); and (4) cognitive bias: CoBBLEr (Koo et al., 2024), against 7 state-of-the-art (SOTA) LLMs, such as GPTs and Claude. We show that our REC models outperform state-of-the-art baselines on these benchmarks, and that the baselines often struggle to provide accurate citation, undermining their reliability.
@@ 2 REC Model Citation Capabilities
We illustrate the content quality citation task in Figure 1, where the auto-evaluator performs content quality evaluation on another LLM’s generation according to a specified metric, and provides rating and explanation with citation for verifiability as an output. When not in an evaluation setting, our models can perform general RAG citation, where we cite from retrieved articles and tag the citation with an answer generated by another LLM, as illustrated in Figure 2.
@@ 3 Related Work
Providing a reference to the attributable source of information (Liu et al., 2023a; Gao et al., 2023) has been proposed as one of the approaches to mitigate hallucination and improve the factuality of LLM generation. Such work fall into the general RAG citation setting defined in our work, where the aim is to generate verifiable content through citation of provided articles. Initial efforts fine-tune LLMs using human-written examples (Nakano et al., 2021) or machine-generated examples verified by humans (Menick et al., 2022), but their privately maintained training data limits further research. With the advent of more capable LLMs (OpenAI, 2023; Jiang et al., 2023a), most existing work rely on zero-shot prompting or few-shot prompting to cite articles during generation (Kamalloo et al., 2023; Gao et al., 2023; Liu et al., 2023a), although the quality of their generated citation leaves much room for improvement (Malaviya et al., 2024). Other work utilize an additional natural language inference (NLI) model to add citation (Gao et al., 2022; Chen et al., 2023). To improve upon prior work, we propose a fine-tuned LLM that can generate better-grounded responses supported with citation of various modes, with the fine-tuning dataset released to the public to facilitate future research and no additional NLI models needed.
General-purpose auto-evaluators
Collecting human annotations to evaluate LLM is not only costly and time consuming, but hard to replicate (Ouyang et al., 2022; Zheng et al., 2023; Chiang & Lee, 2023), and as a result, LLMs become a natural automated proxy to evaluate LLM capabilities on various benchmarks (Bai et al., 2024; Bubeck et al., 2023; Chiang et al., 2023; Fu et al., 2023; Liu et al., 2023b; Wang et al., 2023a). Existing LLM auto-evaluators often judge LLM outputs by expressing “preference” over outputs from a reference model (Dubois et al., 2024; Li et al., 2023b; Yuchen Lin et al., 2024), or by providing rating and explanation according to a user-defined metric as a direct assessment (Li et al., 2023a; Wang et al., 2023b, 2024b). Closer to our work, Jiang et al. (2024) and Xu et al. (2023) fine-tune LLMs to generate rating, explanation, and a detailed analysis to pinpoint errors in a response evaluated. Unlike prior work, our fine-tuned model is a more generalizable auto-evaluator since it provides content quality citation for both good and bad evaluated responses, in addition to rating and explanation. This not only allows users to be able to diagnose where errors are from but provides supporting evidence when a generation is good, which are both important to establish trust in a model.
@@ 4 Data Collection
To fine-tune our general-purpose LLM auto-evaluator, we meticulously collect a mixture of data encompassing a broad spectrum of LLM capabilities (Section 4.1), and we denote our collected dataset as REC-Data. Due to a lack of publicly available content quality citation datasets, we leverage synthetic data generation using an LLM to curate such data (Section 4.2) to facilitate the research community. We perform rule-based automatic quality check followed by a unified task formatting to post-process our data for instruction fine-tuning (Section 4.3).
@@ 4.1 Task Types
To enhance the explainability in the automatic evaluation of our LLM auto-evaluator, while maintaining its general instruction-following capability, we gather datasets from a diverse range of task types (See detailed REC-Data distribution in Appendix C). These tasks represent essential capabilities that an advanced auto-evaluator LLM should possess:
@@ 5 Model Training
Our general-purpose LLM auto-evaluator is available in three sizes, REC-8B, REC-12B, and REC-70B. REC-12B is instruction fine-tuned from Mistral-Nemo [^4] https://mistral.ai/news/mistral-nemo, and REC-8B and REC-70B are instruction fine-tuned from Llama-3.1-8B and Llama-3.1-70B (Team, 2024), respectively. We trained REC-8B and REC-12B to assess the generalizability of the REC approach to different model architectures in similar model sizes. We adopted supervised fine-tuning (SFT) to optimize the models. To accommodate GPU memory constraints: we filtered examples where the combined length of a prompt and a response exceeded 6,144 tokens. We used Low-Rank Adaptation (LoRA) during fine-tuning (Hu et al., 2021), with a rank of $r=256$ for REC-8B and REC-12B, and $r=64$ for REC-70B. We applied 4-bit quantization for REC-70B during initialization. All models were trained using a learning rate of $1\times 10^{-4}$, with data shuffled and for a single epoch. REC-8B and REC-12B were trained with a batch size of 2 and a gradient accumulation factor of 8, whereas for REC-70B we used a batch size of 1 with the same gradient accumulation factor. All models were trained on eight NVIDIA H100 GPUs, each with 80GB of memory. The total training time was approximately 4.5 hours for REC-8B and REC-12B, and 5.5 hours for REC-70B.
@@ 6 Evaluation
In this section, we evaluate our REC models on diverse benchmarks assessing LLMs’ (1) RAG citation capability: ALCE (Gao et al., 2023), and ExpertQA (Malaviya et al., 2024); (2) content quality citation capability: human evaluation on ABCD summarization (Chen et al., 2021); (3) general capabilities: RewardBench (Lambert et al., 2024), and LLM-AggreFact (Tang et al., 2024a); and (4) cognitive bias: CoBBLEr (Koo et al., 2024), each of which is an important measure to a general-purpose LLM auto-evaluator. We compare our results against 7 SOTA LLMs, including Misral-7B (Mistral-7B-Instruct-v0.2(Jiang et al., 2023b), Mistral-Nemo (Mistral-Nemo-Instruct-2407), Llama-3.1-70B (Grattafiori et al., 2024), Claude-3-Opus (claude-3-opus-20240229), GPT-3.5 (gpt-3.5-turbo), GPT-4 (gpt-4-turbo(OpenAI et al., 2024), and GPT-4o (gpt-4o).
@@ 6.1 RAG citation & Correctness
ExpertQA (Malaviya et al., 2024) is designed to evaluate LLMs’ ability to generate accurate and well-attributed responses in technical and high-stakes fields, such as medicine and law. It was developed by involving experts from 32 different fields, who contributed 2,177 domain-specific questions based on their knowledge. In total, 484 participants helped curate these questions. LLMs were then used to generate responses to the questions, followed by human experts evaluating the quality of these responses on several criteria, including factual correctness, completeness of attribution, source reliability, and informativeness. We adopt the metrics used in its paper: AutoAIS (Gao et al., 2022), which is similar to citation recall, and FActscore (Min et al., 2023) which measures the percentage of generated claims that are factual.
@@ 6.2 Rating & Explanation & Citation
To evaluate our LLM auto-evaluator’s citation capability more comprehensively, besides examining its general RAG citation capability on QA tasks as reported in Section 6.1, we evaluate the content quality citation capability on a summarization task, the ABCD dataset (Chen et al., 2021), in this section.
@@
The ABCD dataset (Chen et al., 2021) contains customer support conversation transcripts. We generate summaries by prompting an open-source LLM, Mistral-7B (Jiang et al., 2023a), with: “Summarize knowledge from transcripts after they’ve ended, including the customer issue and resolution.” to form (dialogue, summary) pairs for evaluation. LLM auto-evaluators then take as input the summarization task prompt, and a (dialogue, summary) pair, to generate rating, explanation, and citation according to the metrics defined in Section 1. Note that we use a held-out set of 50 examples in this experiment so there exists no data overlap with the ABCD subsets used during training.
@@
Unlike the general RAG citation benchmarks evaluated in Section 6.1, which contain human annotated citation groundtruth, there exists no well-established benchmarks to evaluate content quality citation. Hence, we seek human annotations to serve as a reference standard in this task. The machine outputs were labeled by 12 machine learning experts, each of whom has a graduate degree in computer science or related fields. They were asked to evaluate the correctness of each of the rating, explanation, and citation generated by the machine. For rating and explanation evaluation, the annotators were asked to provide “Yes/No” labels to determine the correctness of the response. For citation evaluation, they were asked to to provide manually-written citation for the claims in the generated explanation to serve as the groudtruth citation. (Details of the human labeling guideline can be found in Appendix F). Human evaluation on the Coherence metric was not performed because it is conceptually challenging to define what should be cited in a coherent response (e.g., citing an entire paragraph adds little interpretive value). Citation for coherence are more meaningful in negative cases where parts of the response are incoherent. To ensure a balanced evaluation with both positive and negative cases across metrics, we decided to exclude Coherence from the ABCD summarization content quality task.
@@
We assign two labelers for each model and find the average inter-rater agreement is around 95%. We report the average accuracy for rating and explanation across the two sets of labels, and we compute F1 score between machine-generated citation and the intersection of the two sets of manual citation. From Table 3, we see that most LLM auto-evaluators obtain high rating and explanation accuracies, indicating that rating in “Yes/No” and free-form explanation generation are relatively easy tasks for the LLMs. For citation, our REC models perform the best in both completeness and faithfulness metrics, while in instruction-following, they are ranked second to llama-3.1-70B and Claude3-Opus. Specifically, for the instruction-following metric, we find REC models often cite slightly more sentences to ensure a good coverage of the entire conversation between a customer and an agent. In contrast, Llama-70B and Claude3 seem to be more selective.
@@
Table 4: General LLM capabilities evaluation on RewardBench. Scores for all the compared LLMs are obtained from the RewardBench Leaderboard (Lambert et al., 2024).